Find the eigenvalues and eigenvectors of the matrix 6 5 B- [ -5-2
Answer #1

Solution : is The given matrix 5 B 6 5 -2 for to solve we the eigenvalues have characteristics equation, det (B-212) 6-7 5 =>4t3 ( )) (9) - 나 + 3 21 -3i)지+572 G (+3i) 지 -SA + (4+i2 = 0 augmented matrix { i3 The 이 443 5 - 디 +3 4- 31 25 메 T 4-31 。。 5 --4t31 X2 X 5 22 ut X2 C (any Constant Perand not equal to o number then 4 +31 the is X = ☺ (4+0) Ci 5 1 2,-2-31 eigen vectorThe matrix is, augmented 4-3i 5 -5 -4-3, 4+1 RI 25 1 4+31 4+31 0 R2 +5R 1 -4-31 - Г O o The system o is equivalent to, 4+31 XThe eigen eigenvalues and eigen vectors of the matrix are <-4+31 5 for X= eigen vector is 21=2-31 670] for 72 = 2+3; is X = (